OpenStudy (anonymous):

Jenny and Penny can eat a whole gallon of ice cream together in 2 hours. Penny and Lenny can eat a whole gallon of ice cream together in 1 2\3 hours( it is one two by 3 hours) . All three together can eat a whole gallon of ice cream together in 1 hour. How many hours would it take Jenny to eat a whole gallon of ice cream all by herself?

OpenStudy (anonymous):

All I know is that I want to murder whoever thought that rhyming penny,jenny.and lenny was cute and that putting it in a math problem was a good idea -_-

OpenStudy (anonymous):

please lets get this thing done

OpenStudy (anonymous):

do u know how to solve these eqns.?

OpenStudy (anonymous):

yes lall that is given . fine

OpenStudy (anonymous):

sorry i am deleting the above post because i have done wrong..... i would like to tell you how.....

OpenStudy (anonymous):

let time taken by jenny to eat a whole gallon individually =j hr. " " penny " " =p hr. " lenny " = l hr. now amount of ice cream eaten from gallon in 1 hr. by jenny = 1/j " penny = 1/p " lenny = 1/l

OpenStudy (anonymous):

now to make the eqns.

OpenStudy (anonymous):

since jenny and penny can eat the whole one gallon in 2hr. time hence 2 * [(1/j) + (1/p) ] = 1 .................(1)

OpenStudy (anonymous):

do you get the eqn.?

OpenStudy (anonymous):

got to know how it is made?

OpenStudy (anonymous):

Ok..got it now

OpenStudy (anonymous):

see jenny can eat amount ice cream per hr =1/j similarly by penny =1/p but both combines to eat the whole on gallon in 2hr hence you add the amounts of ice cream eaten by penny and jenny in 1 hr but both combines to eat ice cream in one gallon 2hr thus i hv multiplied by 2 to get the amount of one gallon on the rhs of eqn.

OpenStudy (anonymous):

@danyboy9169

OpenStudy (anonymous):

llly other eqns. are (5/3) * [1/p + 1/l ] = 1 ....................(2) 1/j + 1/p + 1/l = 1............................(3)

OpenStudy (anonymous):

now to solve the eqns. from (1) 1/j + 1/p =1/2 .............(4) put (4) in (3) you get 1/2 + 1/l = 1 1/l = 1/2

OpenStudy (anonymous):

llly you can get for jenny but remember in the end you should write t. taken by lenny = l hr. = 2 hr.

OpenStudy (anonymous):

do you get the things ? you can work for jenny...